OpenStudy (anonymous):

find a power series for f(x)=ln(1-x) centered at x=0. someone please help!

OpenStudy (anonymous):

this is a famous one \[-\sum\frac{x^n}{n}\]

OpenStudy (anonymous):

can you show me some work please im lost on this one

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(1-x)=\frac{1}{x-1}=-\frac{1}{1-x}\] and the last one is a well known power series

OpenStudy (anonymous):

\[\frac{1}{1-x}={1+x+x^2+...}=\sum_{k=1}^\infty x^n\] integrate term by term

OpenStudy (anonymous):

so which one is it your first one or somply x^n like your second?

OpenStudy (anonymous):

it is \[\ln(1-x)=-\sum_{k=1}^{\infty}\frac{x^k}{k}\]

OpenStudy (anonymous):

you can also take successive derivatives and see that it works

OpenStudy (anonymous):

ok i get ya. whats the radius of convergance

OpenStudy (anonymous):

\[f(x)=\ln(1-x)\] \[f'(x)=\frac{1}{x-1}=(x-1)^{-1}\] \[f''(x)=-(x-1)^{-2}\] \[f'''(x)=2(x-1)^{-3}\] \[f^{(n)}(x)=(-1)^{n-1}(n-1)!(1-x)^{-n}\]

OpenStudy (anonymous):

plug in 0, divide by \[n!\] and get \[-\frac{x^n}{n}\] for each term

OpenStudy (anonymous):

think this is good on the domain, \[x<1\]